OpenStudy (anonymous):

You decide to enroll in a plan that gives you a 10% discount if you make telephone calls 7pm – 6am. The per minute charge is $0.10 and the sales tax is 6%. On the first month’s bill you used the phone for 130 minutes 6am – 7pm and then 400 minutes 7pm – 6am. How much was your total bill?

OpenStudy (anonymous):

$56.18 $53.00 $51.94 $49.00 ...

OpenStudy (anonymous):

duration of calls made during 7-6= 400 in minutes cost per minute = 0.10 so for 400 min= 400*0.10=40 we also get a 10% discount so cost is = 40*0.10=4 so cost for 7-6 am = 36 dollars this is subject to sales tax of 6 percent so final cost = 0.06*36+36=38.16

OpenStudy (anonymous):

make sure you understand how i obtained 36

OpenStudy (anonymous):

now for 6-7 we have cost = 130*.10=13 with tax of 6% = 13.78

OpenStudy (anonymous):

final cost or bill = 13.78+38.16

OpenStudy (anonymous):

51.94, okay! I see, the only thing I did wrong was I added the sales tax first, so it was more then it really should have been! I understand, thank you!

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

hope you learn it right

OpenStudy (anonymous):

I think I'll get it here soon. Thanks again.

OpenStudy (anonymous):

check the answers thoug

OpenStudy (anonymous):

The two that I just did? That first one we did together was wrong.

OpenStudy (anonymous):

You said it was approx, $51.02, but the answer was $54.75 I guess.

OpenStudy (anonymous):

did we have the data correct lets see

OpenStudy (anonymous):

Do what?

OpenStudy (anonymous):

The answer was- The “Limited Plan” offers a 15% discount for power usage between 7pm – 9am if most of your usage happens during that time. If not, then you lose your discount and you pay a 20% premium on the usage outside of that time. The cost per kilowatt hour is $0.05432. How much would you have paid in March if you used 450 KWH 7pm – 9am and 465 KWH 9am – 7pm? We had all that part correct. but the answer was wrong.

OpenStudy (anonymous):

lets redo it

OpenStudy (anonymous):

Well I got it all described to me when it said I got it wrong, I know how to do it now, thanks.